# Structured Evaluation of Synthetic Tabular Data

### v1

@@ 3 Structured Evaluation Framework
Metric (data type)
Estimate $f_{q},f_{p}$
Score $t$
Implication
@@
TVComplement (categorical)
PMF of $S[c_{j}]$
$1-\mathrm{TV\;dist}(f_{q},f_{p})/Z$
$t=1\Leftrightarrow Q(c_{j})=P(c_{j})$ as $n\rightarrow\infty$
KSComplement (continuous)
CDF of $S[c_{j}]$
$1-\mathrm{KS\;stat}(f_{q},f_{p})$
same as above
Simple-statstics based (either)$\ast$
support, range, mean, median, std of $S[c_{j}]$
$f_{q}/f_{p}$ or $1-|f_{q}-f_{p}|/Z$
$t=1\Leftarrow Q(c_{j})=P(c_{j})$
PCC-Marginal (both)$\star$
empirical CDFs from $\hat{P}(S_{i,j})$
$1-\frac{1}{2}|ppe|$
$t=1\Leftarrow Q(c_{j})=\hat{P}(c_{j})$ as $n\rightarrow\infty$
@@
Contingency-Sim (categorical)
PMF of $S[c_{j},c_{k}]$
$1-$ total variation $\mathrm{dist}(f_{q},f_{p})/Z$
$t=1\Leftrightarrow Q(c_{j},c_{k})=P(c_{j},c_{k})$ as $n\rightarrow\infty$
Correlation-Sim (continuous)
$\rho(S[c_{j}],S[c_{k}])$
$1-|f_{q}-f_{p}|/Z$
$t=1\Leftarrow Q(c_{j},c_{k})=P(c_{j},c_{k})$
MutualInformation-Sim (both)$\star$
$MI(S[c_{j}],S[c_{k}])$
$1-\sum|f_{q}-f_{p}|/Z$
same as above
PCC-Pairwise (both)$\star$
empirical CDFs from $\hat{P}(S_{i,(j,k)})$
$1-\frac{1}{2}|ppe|$
$t=1\Leftarrow Q(c_{j},c_{k})=\hat{P}(c_{j},c_{k})$ as $n\rightarrow\infty$
@@
ML efficacy (categorical)$\ast$
$f_{q}$:$ML(S[c_{j}]\mid S[\{c_{-j}\}])$; $f_{p}$:$X[c_{j}]$
acc between argmax $f_{q}(c_{j}\mid X[\{c_{-j}\}])$ and $X[c_{j}]$
higher $t$ generally implies $Q$ more similar to $P$
Privacy metrics (either)$\ast$
$f_{q}$:$ML(S[\{c_{\alpha}\}]\mid S[\{c_{\beta}\}])$; $f_{p}:X[\{c_{\alpha}\}]$
$1-a(X[\{c_{\alpha}\}];f_{q}(\{c_{\alpha}\}\mid X[\{c_{\beta}\}]))$
higher $t$ generally implies $Q$ less similar to $P$
PCC-LOO (either)$\star$
empirical CDFs from $\hat{P}(S_{i,j}\mid S_{i,\{-j\}})$
$1-\frac{1}{2}|ppe|$
$t=1\Leftarrow Q=\hat{P}$ as $n\rightarrow\infty$
@@
ML-Detection (both)$\ast$
discriminator $f$ of $Q(r)$ trained on $S$
$2(1-ROC\;AUC)$
$t=1\Leftarrow Q=P$ as $n\rightarrow\infty$
PCC-FullJoint (both)$\star$
empirical CDFs from $\hat{P}(S_{i})$
$1-\frac{1}{2}|ppe|$
$t=1\Leftarrow Q=\hat{P}$ as $n\rightarrow\infty$
@@
MissingValue-Sim (either)
fraction of missing values in $S_{\nu}[c_{j}]$
$1-|f_{q}-f_{p}|$
$t=1\Leftrightarrow Q_{\nu}(c_{j})=P_{\nu}(c_{j})$ as $n\rightarrow\infty$
Missing-NotAtRandom-Sim (both)$\star$
correlation matrix formed from column pairs in $S_{\nu}[\{c_{\nu}\}]$
$1-\sum|f_{q}-f_{p}|/Z$
$t=1\Leftarrow Q_{\nu}(\{c_{\nu}\})=P_{\nu}(\{c_{\nu}\})$
Covariate-DependentMissing-Sim (both)$\star$
correlation matrix formed by correlating columns in $S_{\nu}[\{c_{\nu}\}]$ with those in $S_{\nu}[\{c_{-\nu}\}]$
$1-\sum|f_{q}-f_{p}|/Z$
$t=1\Leftarrow Q_{\nu}(\{c_{\nu}\}\mid\{c_{-\nu}\})=P_{\nu}(\{c_{\nu}\}\mid\{c_{-\nu}\})$

### v2

@@ 3 Structured Evaluation Framework
Metric (data type) Estimate $f_{q},f_{p}$ Score $t$ Implication
@@
TVComplement (categorical) PMF of $S[c_{j}]$ $1-\mathrm{TV\;dist}(f_{q},f_{p})/Z$ $t=1\Leftrightarrow Q(c_{j})=P(c_{j})$ as $n\rightarrow\infty$
KSComplement (continuous) CDF of $S[c_{j}]$ $1-\mathrm{KS\;stat}(f_{q},f_{p})$ same as above
Simple-statstics based (either)$\ast$ support, range, mean, median, std of $S[c_{j}]$ $f_{q}/f_{p}$ or $1-|f_{q}-f_{p}|/Z$ $t=1\Leftarrow Q(c_{j})=P(c_{j})$
PCC-Marginal (both)$\star$ empirical CDFs from $\hat{P}(S_{i,j})$ $1-\frac{1}{2}|ppe|$ $t=1\Leftarrow Q(c_{j})=\hat{P}(c_{j})$ as $n\rightarrow\infty$
@@
Contingency-Sim (categorical) PMF of $S[c_{j},c_{k}]$ $1-$ total variation $\mathrm{dist}(f_{q},f_{p})/Z$ $t=1\Leftrightarrow Q(c_{j},c_{k})=P(c_{j},c_{k})$ as $n\rightarrow\infty$
Correlation-Sim (continuous) $\rho(S[c_{j}],S[c_{k}])$ $1-|f_{q}-f_{p}|/Z$ $t=1\Leftarrow Q(c_{j},c_{k})=P(c_{j},c_{k})$
MutualInformation-Sim (both)$\star$ $MI(S[c_{j}],S[c_{k}])$ $1-\sum|f_{q}-f_{p}|/Z$ same as above
PCC-Pairwise (both)$\star$ empirical CDFs from $\hat{P}(S_{i,(j,k)})$ $1-\frac{1}{2}|ppe|$ $t=1\Leftarrow Q(c_{j},c_{k})=\hat{P}(c_{j},c_{k})$ as $n\rightarrow\infty$
@@
ML efficacy (categorical)$\ast$ $f_{q}$:$ML(S[c_{j}]\mid S[\{c_{-j}\}])$; $f_{p}$:$X[c_{j}]$ acc between argmax $f_{q}(c_{j}\mid X[\{c_{-j}\}])$ and $X[c_{j}]$ higher $t$ generally implies $Q$ more similar to $P$
Privacy metrics (either)$\ast$ $f_{q}$:$ML(S[\{c_{\alpha}\}]\mid S[\{c_{\beta}\}])$; $f_{p}:X[\{c_{\alpha}\}]$ $1-a(X[\{c_{\alpha}\}];f_{q}(\{c_{\alpha}\}\mid X[\{c_{\beta}\}]))$ higher $t$ generally implies $Q$ less similar to $P$
PCC-LOO (either)$\star$ empirical CDFs from $\hat{P}(S_{i,j}\mid S_{i,\{-j\}})$ $1-\frac{1}{2}|ppe|$ $t=1\Leftarrow Q=\hat{P}$ as $n\rightarrow\infty$
@@
ML-Detection (both)$\ast$ discriminator $f$ of $Q(r)$ trained on $S$ $2(1-ROC\;AUC)$ $t=1\Leftarrow Q=P$ as $n\rightarrow\infty$
PCC-FullJoint (both)$\star$ empirical CDFs from $\hat{P}(S_{i})$ $1-\frac{1}{2}|ppe|$ $t=1\Leftarrow Q=\hat{P}$ as $n\rightarrow\infty$
@@
MissingValue-Sim (either) fraction of missing values in $S_{\nu}[c_{j}]$ $1-|f_{q}-f_{p}|$ $t=1\Leftrightarrow Q_{\nu}(c_{j})=P_{\nu}(c_{j})$ as $n\rightarrow\infty$
Missing-NotAtRandom-Sim (both)$\star$ correlation matrix formed from column pairs in $S_{\nu}[\{c_{\nu}\}]$ $1-\sum|f_{q}-f_{p}|/Z$ $t=1\Leftarrow Q_{\nu}(\{c_{\nu}\})=P_{\nu}(\{c_{\nu}\})$
Covariate-DependentMissing-Sim (both)$\star$ correlation matrix formed by correlating columns in $S_{\nu}[\{c_{\nu}\}]$ with those in $S_{\nu}[\{c_{-\nu}\}]$ $1-\sum|f_{q}-f_{p}|/Z$ $t=1\Leftarrow Q_{\nu}(\{c_{\nu}\}\mid\{c_{-\nu}\})=P_{\nu}(\{c_{\nu}\}\mid\{c_{-\nu}\})$